Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax2 + bx + c = 0.
Solutions: -9 and 6.


Answer by stanbon(75887)

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax2 + bx + c = 0.
Solutions: -9 and 6.
-----
y = (x+9)(x-6)
----
y = x^2 + 3x - 54
==================================
Cheers,
Stan H.